OpenStudy (anonymous):

HELP PLEASE compute the modulus and arguement a.1+i b.4-4i c.2+5i d.2(cos(2pi/3)+i sin(2pi/3))

sam (.sam.):

1+i Modulus \[|z|=\sqrt{1^2+1^2}\] \[|z|=\sqrt{2}\]

sam (.sam.):

Argument \[\theta=\tan^{-1}\frac{1}{1}=\frac{\pi}{4}\]

sam (.sam.):